OpenStudy (anonymous):

I just had my math final and I got stuck with this problem. What is the answer? 3 = -1 (x+11) + 1 4 4x 2x Lower numbers are the denominators

OpenStudy (anonymous):

you were supposed to find x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Common denominator

OpenStudy (anonymous):

8 or 4?

OpenStudy (anonymous):

(-1(x+11)+2 ) /4x (-x-11+2)/4x 3/4=(-x-9)/4x Time both side by 4x 12x/4=-x-9 3x=-x-9 4x=-9 x=-9/2 if I didn't screw anything up

OpenStudy (anonymous):

well thanks but it already too late i got it wrong on my final :(

OpenStudy (amistre64):

multiply it all by 4x to clear the fractions

OpenStudy (amistre64):

3 = -1 (x+11) + 1 ---------------- * 4x 4 4x 2x 3x = -1 (x+11) + 2

OpenStudy (amistre64):

3x = -x -11+2 4x = -9 x = -9/4

OpenStudy (anonymous):

Thanks guys - still bummin'

OpenStudy (amistre64):

oh dont worry, you got plenty of time to be right in the future :)

OpenStudy (amistre64):

i gotta wait a few more hours to see if i read my statistics test right; I dont recall solving a problem according to the question. but there are lots of things I dont remember these days soo.... it mighta been right lo

OpenStudy (anonymous):

good luck with that